Some recent financial statements for Smolira Golf Corp. follow GOLF C 2014 and 2015 Balance Sheets Liabilities and Owners Equity 2014 Assets 2015 2014 2015 Current assets Current liabilities Accounts payable Notes payable Other $ 24,284 13,000 12,671 28,200 11,900 19,900 $ 23,056 13,548 26,982 25,200 16,300 28,200 Accounts receivable 49,955 60,000 $ 80,000 92,000 Total S 63,586 69,700 Total Long-term debt Owners equity Common stock and paid-in surplus Accumulated retained eamings 60,000 $ 60,000 222,000 209,326 Fixed assets Net plant and equipment 269,326 $282,000 $399,281 $434,000 $335,695 $364,300 Total Total assets $399,281 $434,000 Total liabilities and owners equity SMOLIRA GOLF CORP 2015 Income Statement Sales Cost of goods sold Depreciation 375,123 258,000 45,600 Earnings before interest and taxes Interest paid s 71,523 15,400 Taxable income Taxes (40%) s 56,123 22,449 Net income $33,674 $21,000 12,674 Retained earnings
media%2Fe9f%2Fe9fc1fb6-f414-4479-863e-19
Answer #1

ROE = (Net Income/Sales) X (Sales/Average Assets) X (Average Assets/Avenge Equity)

PROFIT MARGIN = 33674/375123 =0.0898 = 8.98%

TOTAL ASSETS TURNOVER = 375123/(399281 + 434000/2) = 0.90

EQUITY MULTIPLIER = (399281 + 434000/2)/ (269326 + 2820000)/2 = 1.51

ROE = 8.98% X 0.90 X 1.51 = 12.20%

Go through it, Any doubts, please feel free to ask, Give positive feedback, Thank you